Problem 10-9 (Algo) Comprehensive Variance Analysis [LO10-1, LO10-2, LO10-3]

Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 980 hours each month to produce 1,960 sets of covers. The standard costs associated with this level of production are:

Total Per Set
of Covers
Direct materials $ 32,340 $ 16.50
Direct labor $ 6,860 3.50
Variable manufacturing overhead (based on direct labor-hours) $ 1,960 1.00
$ 21.00

During August, the factory worked only 1,000 direct labor-hours and produced 2,100 sets of covers. The following actual costs were recorded during the month:

Total Per Set
of Covers
Direct materials (6,000 yards) $ 34,020 $ 16.20
Direct labor $ 7,770 3.70
Variable manufacturing overhead $ 3,990 1.90
$ 21.80

At standard, each set of covers should require 2.5 yards of material. All of the materials purchased during the month were used in production.

Required:

1. Compute the materials price and quantity variances for August.

2. Compute the labor rate and efficiency variances for August.

3. Compute the variable overhead rate and efficiency variances for August.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

Answer:

Standard materials price per yard 6.6 =16.50/2.5
Standard labor hour per unit 0.5 =980/1960
Standard labor rate per labor hour 7 =3.5/0.5
Standard variable rate per labor hour 2 =1/0.5
1
Materials price variance 5580 F =34020-(6000*6.6)
Materials quantity variance 8910 U =6.6*(6000-2100*3.5)
2
Labor rate variance 770 U =7770-(1000*7)
Labor efficiency variance 350 F =7*(1000-2100*0.5)
3
Variable overhead rate variance 1990 U =3990-(1000*2)
Variable overhead efficiency variance 100 F =2*(1000-2100*0.5)